# Collaborative Feature-Logits Contrastive Learning for Open-Set Semi-Supervised Object Detection

### v1

@@ 2.2. Semi-Supervised Object Detection
Semi-supervised learning enhances model performance by leveraging labeled and unlabeled data, reducing reliance on fully labeled datasets compared to weakly supervised approaches (Zhang et al., 2023b; Jiang et al., 2021). Recent advancements in semi-supervised object detection (SSOD) have shown substantial promise, particularly through methods such as data augmentation (Wang et al., 2021; Li et al., 2022a), consistency regularization (Berthelot et al., 2019; Jeong et al., 2019; Zheng et al., 2022), and pseudo-labeling techniques (Zhou et al., 2021; Li et al., 2022b; Liu et al., 2021). Consistency-based methods enforce consistency constraints to utilize unlabeled data, while pseudo-label-based methods generate pseudo-labels to guide training. However, these methods are limited to close-set scenarios, as the presence of OOD class can degrade pseudo-label quality, leading to semantic confusion and reduced performance.

### v2

@@ 2.2. Semi-Supervised Object Detection
Semi-supervised learning enhances model performance by leveraging labeled and unlabeled data, reducing reliance on fully labeled datasets compared to weakly supervised approaches (Zhang et al., 2023c; Jiang et al., 2021). Recent advancements in semi-supervised object detection (SSOD) have shown substantial promise, particularly through methods such as data augmentation (Wang et al., 2021; Li et al., 2022a), consistency regularization (Berthelot et al., 2019; Jeong et al., 2019; Zheng et al., 2022; Zhang et al., 2023b), and pseudo-labeling techniques (Zhou et al., 2021; Li et al., 2022b; Liu et al., 2021). Consistency-based methods enforce consistency constraints to utilize unlabeled data, while pseudo-label-based methods generate pseudo-labels to guide training. However, these methods are limited to close-set scenarios, as the presence of OOD class can degrade pseudo-label quality, leading to semantic confusion and reduced performance.